A battery of emf 12v and internal resistance 1.5 ohms connects across a 10 ohm resistor.

Calculate the current of the circuit

Calculate the terminal voltage


1
Expert's answer
2021-02-24T12:45:06-0500

Calculate effective resistance,

"R_E=1.5\Omega +10\Omega=11.5\Omega"

"I=\frac{V}{R}=\frac{12}{11.5}=1.0435A"

Terminal voltage"=IR=1.0435\times 1.5=1.5653v"